A 5.50 L flask of CH₄ at 35.0°C and 900. torr is heated to 75.0°C. What is the new pressure?


1
Expert's answer
2021-03-05T03:29:11-0500

According to Gay-lussac's law

Increase in temperature produces increase in pressure

P/T = K

P1 = 900 torr = (900/760)atm = 1.18atm

T1 = 35°C = (273 + 35)K = 308K

P2 = ?

T2 = 75°C = (273 + 75) = 248K

P2 = (P1 x T2)/ T1

= (1.18 x 348)/ 308

= 1.33atm = 1010.80 torr

Hence, the new pressure is 1010.80 torr.

This proves that Gay-lussac is right